What is the x-coordinate of the vertex of the parabola whose equation is y = 3x2 + 12x + 5?
x = -4 , x= - 3 , OR x= - 2

Answer by stanbon(75887)

What is the x-coordinate of the vertex of the parabola whose equation is y = 3x2 + 12x + 5?
a = 3 ; b = 12 ; c = 5
---
Vertex occurs where x = -b/2a = -12/(2*3) = -2
----------------------------
Cheers,
Stan H.
==========================

x = -4 , x= - 3 , OR x= - 2